OpenStudy (anonymous):

Suppose y is inversely proportional to x and the constant of proportionality equals 4. What is the minimum of the sum of x and y if x and y are both positive?

OpenStudy (lgbasallote):

y = 4/x huh? minimum...does this have something to do with calculus?

OpenStudy (anonymous):

yes this is a calculus problem

OpenStudy (bahrom7893):

y=4x^(-1) y'=-4/x^2

OpenStudy (bahrom7893):

ohh wait lol wrong approach

OpenStudy (bahrom7893):

minimize x+y

OpenStudy (bahrom7893):

x+y=x+4x^(-1) minimize f(x)=x+4x^(-1)

OpenStudy (bahrom7893):

f'(x)= 1 -4x^(-2)

OpenStudy (anonymous):

so take derivative and set it to zero?

OpenStudy (bahrom7893):

1-4x^(-2)=0 x^2-4=0 x=+/-2 but since x is positive, we take x=2

OpenStudy (bahrom7893):

that means y=4/x=4/2=2

OpenStudy (anonymous):

the answer's 2? i tried 2 earlier but my hw site said it was wrong.

OpenStudy (bahrom7893):

minimum is 4.. 2+2 is 4

OpenStudy (anonymous):

o

OpenStudy (anonymous):

wait how did you go from that 1-4x^(-2)=0 to that x^2-4=0